# Higher Order Dualities between Prime Ideals

### v1

@@ 5 Two Estimates of Sums involving $\mu_{K}(I)$ and $\omega_{K}(I)$
$\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{T}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)\sum_{T<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})\ll\sum_{N(J)\leq\frac{X}{T}}\frac{X}{N(J^{\prime})e^{c_{6}\left(\log\frac{X}{N(J^{\prime})}\right)^{\frac{1}{12}}}}\ll\frac{X\log X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}$ (5.13)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)\ll\frac{X\log X}{e^{c_{8}(\log X)^{\frac{1}{12}}}}$ (5.14)
Finally, combining (5.9), (5.11), (5.12), (5.13) and (5.14), we have
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\ll\frac{X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log T(\log\log X)^{2}}{\log\left(\frac{X}{T}\right)}+\frac{X\log X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log X}{e^{c_{8}(\log X)^{\frac{1}{12}}}}$ (5.15)
@@
Both the above theorems are important applications of our duality identities. Not only does the duality help us get a new formula for the Chebotarev Density, it also helps us estimate different sums over restricted sets of ideals. Higher order dualities can similarly be used to deduce such estimates of other sums.131313one can follow methods in [AS(ip)] to get estimates of new sums
@@ 6 Duality for Arbitrary Sets of Prime Ideals
In §2, we prove a general higher order duality identity (Theorem 2.4) involving prime ideals of $k^{th}$ largest norm and the smallest norm. But the duality has been specific to only a particular condition, i.e. for prime ideals being unramified in the bigger field of extension and satisfying the Artin Symbol condition for a fixed conjugacy class $C$ of a Galois group of a corresponding Galois extension. But we can extend the duality identity for a higher class of conditions without having to hardly change anything in the proofs. In this section, we will state141414the proof will be omitted as they exactly follow the same idea as Theorem 2.3 and Theorem 2.4 the general duality for arbitrary but fixed sets of prime ideals in a given ring of integers.
@@
$S(\mathcal{A})$ $:=\{I\subset\mathcal{O}_{K}:I\text{ is salient with unique }\mathfrak{p};\;\mathfrak{p}\in\mathcal{A}\}$
$Q_{\mathcal{A}}^{k}(I)$ $:=\#\{\mathfrak{p}\subset\mathcal{O}_{K}:\mathfrak{p}\;\text{is a prime};\;I\subset\mathfrak{p};\;M_{k}(I)=N(\mathfrak{p});\;\mathfrak{p}\in\mathcal{A}\}$
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=(-1)^{k}Q_{\mathcal{A}}^{k}(I)$
@@
We can now use Theorem 6.1 to obtain density type results, similar to our Theorem 4.3 above. We will, in our hypothesis, assume that the averages of the sets $Q_{\mathcal{A}}^{k}(I)$ holds. We state such a theorem for $k=2$ below.
@@
$\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{1}(I)\sim\mathcal{K}\cdot X\quad and\quad\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{2}(I)\sim\mathcal{K}\cdot X$
@@
$\lim_{x\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\mathcal{K}$
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1+O\left(\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right)$ (6.1)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\ll\frac{X\log\log X}{e^{c_{6}(\log X)^{\frac{1}{12}}}}$ (6.2)
@@
$=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ (6.3)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)$ $=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(J)\leq X\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)=\frac{1}{c_{K}}\cdot\mathcal{K}\cdot X+o(X)$ (6.4)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ $=-\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}Q_{C}^{1}(I)=-\frac{1}{c_{K}}\cdot\mathcal{K}\cdot X+o(X)$ (6.5)
@@
$X\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=o(X)\implies\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=o(1)$ (6.6)
@@
Choosing the set $\mathcal{A}$ to be the prime ideals that are unramified in $L$ such that their Artin symbols are all $C$, where $L/K$ is a Galois extension and $C\subset G=Gal(L/K)$, we see that $\mathcal{K}=c_{K}\cdot\frac{|C|}{|G|}$, and hence, Theorem 4.3 follows qualitatively from Theorem 6.3.
It is always a challenge to prove the asymptotics in the hypothesis of Theorem 6.3. In our case, Theorem 4.1 and Theorem 1.1 in [SW18] are extremely important results and Theorem 4.3 is a consequence of both of these results. It will be worthwhile to study such collections $\mathcal{A}$ of prime ideals in given rings of integers that satisfy such asymptotics and how the constant $\mathcal{K}$ differs in each case. Of course, if the constants are different, say $\mathcal{K}_{1},\mathcal{K}_{2}$ respectively for $k=1,2$, we will have
$\lim_{x\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=\frac{1}{c_{K}}(\mathcal{K}_{2}-\mathcal{K}_{1})$
There have been works by Kural, McDonald and Sah [KMS20] and Wang [Wa21] that extends works of Alladi, Dawsey, Sweeting and Woo. It is hoped that this general duality and the new approach towards attaining formulas of the Chebotarev Density can now open up even more generalized extensions of results that have been already proved by the authors mentioned above. Sahoo and Jha [JS(ip)] have also worked out the function field analogue of the Alladi duality. A function field analogue of this higher order duality between prime ideals will lead to a complete duality identity for global fields which might even lead to further exploration.
Acknowledgements: I am indebted to my Doctoral Advisor, Prof. K. Alladi for his support and guidance throughout this work. I also sincerely thank Prof. A. Vatwani, and her PhD students Mr. J. Sahoo and Mr. P. N. Jha for helping me out with new explanations and valuable insights that led to the completion of this work.

### v2

@@ 5 Two Estimates of Sums involving $\mu_{K}(I)$ and $\omega_{K}(I)$
$\sum_{\begin{subarray}{c}N(J)\leq\frac{X}{T}\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)\sum_{T<N(J^{\prime})\leq\frac{X}{N(J)}}\mu_{K}(J^{\prime})\ll\sum_{N(J)\leq\frac{X}{T}}\frac{X}{N(J^{\prime})e^{c_{6}\left(\log\frac{X}{N(J^{\prime})}\right)^{\frac{1}{12}}}}\ll\frac{X\log X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}.$ (5.13)
@@
$\sum_{N(I)\leq X}\mu_{K}(I)f(I)\ll\frac{X\log X}{e^{c_{8}(\log X)^{\frac{1}{12}}}}.$ (5.14)
Finally, combining (5.9), (5.11), (5.12), (5.13), and (5.14), we have
$\sum_{\begin{subarray}{c}N(I)\leq X\\ I\in S(L/K;C)\end{subarray}}\mu_{K}(I)\omega_{K}(I)\ll\frac{X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log T(\log\log X)^{2}}{\log\left(\frac{X}{T}\right)}+\frac{X\log X}{e^{c_{6}(\log T)^{\frac{1}{12}}}}+\frac{X\log X}{e^{c_{8}(\log X)^{\frac{1}{12}}}}.$ (5.15)
@@
Both the above theorems are important applications of our duality identities. Not only does the duality help us get a new formula for the Chebotarev Density, but it also helps us estimate different sums over restricted sets of ideals. Higher order dualities can similarly be used to deduce such estimates of other sums.131313one can follow methods in [AS(ip)] to get estimates of new sums
@@ 6 Duality for Arbitrary Sets of Prime Ideals
In §2, we prove a general higher order duality identity (Theorem 2.4) involving prime ideals of $k^{th}$ largest norm and the smallest norm. But the duality has been specific to only a particular condition, i.e., for prime ideals being unramified in the bigger field of extension and satisfying the Artin Symbol condition for a fixed conjugacy class $C$ of a Galois group of a corresponding Galois extension. But we can extend the duality identity for a higher class of conditions without having to hardly change anything in the proofs. In this section, we will state141414the proof will be omitted as they exactly follow the same idea as Theorem 2.3 and Theorem 2.4 the general duality for arbitrary but fixed sets of prime ideals in a given ring of integers.
@@
$S(\mathcal{A})$ $:=\{I\subset\mathcal{O}_{K}:I\text{ is salient with unique }\mathfrak{p};\;\mathfrak{p}\in\mathcal{A}\},$
$Q_{\mathcal{A}}^{k}(I)$ $:=\#\{\mathfrak{p}\subset\mathcal{O}_{K}:\mathfrak{p}\;\text{is a prime};\;I\subset\mathfrak{p};\;M_{k}(I)=N(\mathfrak{p});\;\mathfrak{p}\in\mathcal{A}\}.$
@@
$\sum_{J\supset I}\mu_{K}(J){{\omega_{K}(J)-1}\choose k-1}f(J)=(-1)^{k}Q_{\mathcal{A}}^{k}(I).$
@@
We can now use Theorem 6.1 to obtain density type results, similar to our Theorem 4.3 above. We will, in our hypothesis, assume that the averages of the sets $Q_{\mathcal{A}}^{k}(I)$ exist. We state such a theorem for $k=2$ below.
@@
$\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{1}(I)\sim\mathcal{K}\cdot X\quad and\quad\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{2}(I)\sim\mathcal{K}\cdot X,$
@@
$\lim_{x\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=0.$
@@
$=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\sum_{\begin{subarray}{c}J\subset\mathcal{O}_{K}\\ N(J)\leq X\\ I\supset J\end{subarray}}1+O\left(\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\right).$ (6.1)
@@
$\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\mu_{K}(I)\omega_{K}(I)\left(\frac{X}{N(I)}\right)^{1-\frac{1}{d}}\ll\frac{X\log\log X}{e^{c_{6}(\log X)^{\frac{1}{12}}}}.$ (6.2)
@@
$=\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)+\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I).$ (6.3)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)(\omega_{K}(I)-1)f(I)$ $=\frac{1}{c_{K}}\sum_{\begin{subarray}{c}2\leq N(J)\leq X\\ Q^{1}(J)=1\end{subarray}}Q_{C}^{2}(J)=\frac{1}{c_{K}}\cdot\mathcal{K}\cdot X+o(X),$ (6.4)
@@
$\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}\sum_{\begin{subarray}{c}I\supset J\end{subarray}}\mu_{K}(I)f(I)$ $=-\frac{1}{c_{K}}\sum_{2\leq N(J)\leq X}Q_{C}^{1}(I)=-\frac{1}{c_{K}}\cdot\mathcal{K}\cdot X+o(X).$ (6.5)
@@
$X\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=o(X)\implies\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=o(1).$ (6.6)
@@
We can prove the following corollary too in the similar fashion as of Theorem 6.3.
Corollary 6.3.1.
$\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{1}(I)\sim\mathcal{K}\cdot X\quad and\quad\sum_{2\leq N(I)\leq X}Q_{\mathcal{A}}^{2}(I)\sim\mathcal{K}\cdot X,$
then
$\lim_{x\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)(\omega_{K}(I)-1)}{N(I)}=\frac{1}{c_{K}}\cdot\mathcal{K}.$
@@
Choosing the set $\mathcal{A}$ to be the prime ideals that are unramified in $L$ such that their Artin symbols are all $C$, where $L/K$ is a Galois extension and $C\subset G=Gal(L/K)$, we see that $\mathcal{K}=c_{K}\cdot\frac{|C|}{|G|}$, and hence, Theorem 4.3 follows qualitatively from Corollary 6.3.1.
It is always a challenge to prove the asymptotics in the hypothesis of Theorem 6.3. In our case, Theorem 4.1 and Theorem 1.1 in [SW18] are extremely important results, and Theorem 4.3 is a consequence of both of these results. It will be worthwhile to study such collections $\mathcal{A}$ of prime ideals in given rings of integers that satisfy such asymptotics and how the constant $\mathcal{K}$ differs in each case. Of course, if the constants are different, say $\mathcal{K}_{1},\mathcal{K}_{2}$ respectively for $k=1,2$, we will have
$\lim_{x\to\infty}\sum_{\begin{subarray}{c}2\leq N(I)\leq X\\ I\in S(\mathcal{A})\end{subarray}}\frac{\mu_{K}(I)\omega_{K}(I)}{N(I)}=\frac{1}{c_{K}}(\mathcal{K}_{2}-\mathcal{K}_{1}).$
There have been works by Kural, McDonald, and Sah [KMS20] and Wang [Wa21] that extends works of Alladi, Dawsey, Sweeting, and Woo. It is hoped that this general duality and the new approach to obtaining formulas for the Chebotarev Density can now open up even more general extensions of results already proved by the authors mentioned above. Sahoo and Jha [JS(ip)] have also worked out the function field analogue of the Alladi duality. A function field analogue of this higher order duality between prime ideals will lead to a complete duality identity for global fields, which might even lead to further exploration.
Acknowledgements: I am indebted to my Doctoral Advisor, Prof. K. Alladi, for his support and guidance throughout this work. I also sincerely thank Prof. A. Vatwani and her PhD students, Mr. J. Sahoo and Mr. P. N. Jha, for helping me out with new explanations and valuable insights that led to the completion of this work.